OpenStudy (anonymous):

What is the side length of a cube with a volume of 2197 cm3? @Pawtpie

OpenStudy (anonymous):

the cube root of 2197 is the answer

OpenStudy (anonymous):

Okay so now we need the cube root of that large number of 2197

OpenStudy (anonymous):

^

OpenStudy (anonymous):

Try a random number like 18 cubed first

OpenStudy (anonymous):

18*18*18

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

5832

OpenStudy (anonymous):

Ok that's way too high then, what about 14

OpenStudy (anonymous):

cubed

OpenStudy (anonymous):

2744

OpenStudy (anonymous):

Oh close, 13?

OpenStudy (anonymous):

I'm sure there's a better way to solve these

OpenStudy (anonymous):

2197

OpenStudy (anonymous):

use a calculator

OpenStudy (anonymous):

next right

sammixboo (sammixboo):

Here is the formula to find the volume of a cube \(\rm s^3~=~V\) Well, we know the Volume (V) = 2197 So plug in the value given \(\rm s^3~=~2197\) Do you know how to solve this?

sammixboo (sammixboo):

s = side

OpenStudy (anonymous):

@Pawtpie

sammixboo (sammixboo):

Well, \(\rm s^3~=~2197\) \(\rightarrow\) \(\rm s~=~2197^{(\dfrac{1}{3})}\)

OpenStudy (anonymous):

13

sammixboo (sammixboo):

Correct =) So \(\rm 13^3~=~2197\) \(\rightarrow\) \(\rm 13\times13\times13~=~2197\) \(\rightarrow\) \(\rm 2197~=~2197\)